# Vision Language Modeling of Content, Distortion and Appearance for Image Quality Assessment

### v2

@@ IV-B Results and Comparisons
We test our model on the above mentioned 8 IQA datasets and compare it with 15 BIQA models. The competitors include 16 BIQA methods: BRISQUE [24], NIQE[25], CORNIA [48], WaDIQaM [2], PQR [51], DB-CNN[52], P2P-BM[49], HyperIQA[39], MetaIQA[62], MUSIQ[17], UNIQUE [53], CONTRIQUE[22], Re-IQA[34], SAWAR [60], CLIP-IQA[44], and GRepQ[38]. All results of BIQA are either from the original papers or reproduced by their source codes.
Two widely-used criteria Spearman rank-order correlation coefficient (SROCC) and the Pearson linear correlation coefficient (PLCC) are used to measure the performance of different BIQA models. Table I and Table II list the quantitative results on synthetic distortion datasets and authentic distortion datasets, respectively. The best and second-best results are highlighted in bold. The proposed method achieves 16 best results across all IQA datasets. It indicates that our model exhibits excellent performances on both synthetic and authentic distortions. We attribute it to the data preparation of both synthetic distortion and authentic images for the joint contrast learning. In contrast, only synthetic distortion images can be employed to train the auto-encoders of SAWAR in a self-supervised way. Thus, the performance of SAWAR on authentic distortion datasets is inferior, although its performance on synthetic distortion datasets is impressive. In the CONTRIQUE and Re-IQA models, both types of distortions can be employed to train the quality-relevant feature extractors. However, they only consider visual contrast learning and ignore the benefits of language supervised learning. In addition, the CLIP-IQA model is mainly devoted to exploit the knowledge already contained in the pre-trained CLIP and finetune with simple statements. Different from them, the proposed method based on vision-language and visual contrastive learning can achieve superior results by using a more comprehensive text description.
IV-C T-SNE Visualization of Quality Relevant Features
The success of SLIQUE relies on its image encoder’s ability to extract discriminative image features that can capture information related to factors affecting image quality such as semantic content, distortion characteristics and appearance properties. Fig. 3 shows examples of t-SNE visualization of the learned representations of SLIQUE and that of a recent method CONTRIQUE [22]. We used 750 images sampled from CSIQ (Synthetic distortions - #600) and User-Generated Content (UGC) images from KonIQ (UGC -#150) and pristine images from KADIS (Natural images - #150). It is clearly seen that, for the CONTRIQUE model, some features from differently distorted images are not well distinguished from each other, e.g., the JPEG and the Pink noise. In contrast, the representations learned by our new SLIQUE can better separate different distortion types which may explain why it has superior performances.
@@ IV-C T-SNE Visualization of Quality Relevant Features
Figure 3: 2D t-SNE visualizations of learned representations. For both CONTRIQUE and SLIQUE, we conducted 2D t-SNE visualization experiments using 7 different types of distorted images from 3 image databases.
IV-D Cross Dataset Evaluation
To show the generalization ability of SLIQUE, we further conduct cross dataset evaluations by training and testing on different datasets. Specifically, the ridge regression of SLIQUE is learned by all the data from a dataset while the parameters of image encoder $F$ are still frozen. Then, the model is tested on all the data from another dataset. Four datasets, including two synthetic and two authentic distortion datasets, are involved in the cross dataset evaluation. The results based on SROCC are presented in Table III, where the performances of two recently-developed models are also listed. As shown in Table III, our model outperforms the Re-IQA and the SAWAR models, although all these three models take advantage of unsupervised learning to obtain quality-relevant features. It demonstrates that our joint contrastive learning is highly effective in representing the image quality.
TABLE III: Cross database evaluations of IQA models.
@@ IV-E Ablation Studies
Individual contrastive learning. We first conducted an ablation study to validate the importance of joint contrastive learning by disabling one of the branches. When disabling the image-image branch, only image-language pairs are used to train the encoders $G$ and $F$ from scratch. Similarly, if the image-language branch is disabled, only image-image pairs are employed to train the image encoder $F$. The results based on SROCC are given in Table IV, which shows a performance reduction if we disable either branch. Moreover, we can see from the results that the visual contrastive learning is greatly important, and the vision-language contrastive learning also provides essential benefits.
TABLE IV: Performance of individual contrastive learning.
@@
Starting point of optimization. In the joint training of image-language branch and image-image branch, we optimize the language encoder $G$ and the image encoder $F$ from scratch. An alternative strategy is to optimize them (or one of them) from pre-trained models. Here we employed two well-known pre-trained models, i.e., CLIP and SIMCLR, as the starting point of optimization. The results on synthetic distortion datasets and authentic distortion datasets are recorded in Table V and Table VI, respectively. In the first row of each table, the language encoder $G$ is initialized by the pretrained CLIP weights, and the image encoder $F$ is initialized by the pretrained SIMCLR weights. Other cases are included in the following rows. As the SIMCLR only contains image encoder, it can not used to initialize $G$. From the results in Table V and Table VI, we find that the optimizations from pretrained models do not result in the best performances, although some performances from them are comparable with state of the art. It may be because that they are originally trained in object recognition or image classification tasks. Thus, using pretrained models is not optimal for the image quality assessment task.
TABLE V: Optimization from pretrained models versus from scratch (synthetic distortion datasets).
@@
TABLE VI: Optimization from pretrained models versus from scratch (authentic distortion datasets).
@@
Components in annotation texts. To investigate the impact of different components in the annotation texts on quality prediction, we conducted ablation studies by excluding some components in the textual descriptions, resulting in 4 cases. The results based on SROCC are shown in Table VII. In the first row of Table VII, we only keep the content annotation texts, and exclude both the distortion texts for synthetic distortion images and the appearance texts for authentic distortion images. The results in the first, second, and third rows respectively indicate that using only content text, distortion text, or appearance attribute text alone is insufficient. In the fourth row, we add the distortion information of synthetic distortion images to the textual descriptions. The results show that the performance on authentic distortion datasets KonIQ and CLIVE even decreases, although the performance on synthetic distortion datasets is improved. It may be attributed to the large differences between the synthetic distortion and the authentic distortion. Similarly, the results of adding the appearance texts are provided in the fifth row which is worse than the full model. In the sixth row, we report the results of excluding content texts, demonstrating that the content texts are also beneficial to the IQA task.
TABLE VII: Ablation studies on components in annotation texts. In each column, the top performance is boldfaced.
@@
TABLE VIII: SROCC performances of our model for different amount of training data.
@@
Usage of language feature. In our SLIQUE, only the image feature is employed in the prediction of visual quality of images. An alternative strategy is to use an image captioning method to produce a textual description for the image to be accessed. With such a strategy, the feature from language encoder can also be fed to the ridge regression. Here we employed the captioning method in [26] to produce annotation texts. The results on synthetic distortion datasets and authentic distortion datasets are recorded in Table IX and Table X, respectively. From the results, we can see that using the language feature is not helpful. It is mainly because the texts produced from the captioning method contain little distortion and appearance information. Furthermore, even if the captioning method can produce perfect texts for the IQA task, the language feature would be redundant due to its similarity to the corresponding image feature.
TABLE IX: Exploring more features to use in evaluation of synthetic distortion datasets.
@@
TABLE X: Exploring more features used in the evaluation of authentic distortion datasets.

### v3

@@ IV-B Results and Comparisons
We test our model on the above mentioned 8 IQA datasets and compare it with 15 BIQA models. The competitors include 16 BIQA methods: BRISQUE [57], NIQE[7], CORNIA [58], WaDIQaM [10], PQR [59], DB-CNN[12], P2P-BM[8], HyperIQA[5], MetaIQA[11], MUSIQ[9], UNIQUE [60], CONTRIQUE[31], Re-IQA[32], SAWAR [30], CLIP-IQA[27], and GRepQ[17]. Additionally, we evaluate our model on other types of IQA datasets and compare it with models including HyperIQA[5], MUSIQ[9], CONTRIQUE[31], CLIP-IQA[27] on underwater images UIF[61], UID2021[62], super resolution images QADS[63] and human facial images GFIQA[64]. All results of BIQA are either from the original papers or reproduced by their source codes.
Two widely-used criteria Spearman rank-order correlation coefficient (SROCC) and the Pearson linear correlation coefficient (PLCC) are used to measure the performance of different BIQA models. Table I and Table II list the quantitative results on synthetic distortion datasets and authentic distortion datasets, respectively. The best and second-best performances are highlighted in bold. The proposed method achieves 16 best results across all datasets. It indicates that our model exhibits excellent performances on both synthetic and authentic distortions. We attribute it to the data preparation of both synthetic distortion and authentic images for the joint contrast learning. In contrast, only synthetic distortion images can be employed to train the auto-encoders of SAWAR in a self-supervised way. Thus, the performance of SAWAR on authentic distortion datasets is inferior, although its performance on synthetic distortion datasets is impressive. In the CONTRIQUE and Re-IQA models, both types of distortions can be employed to train the quality-relevant feature extractors. However, they only consider visual contrast learning and ignore the benefits of language supervised learning. In addition, the CLIP-IQA model is mainly devoted to exploit the knowledge already contained in the pre-trained CLIP and finetune with simple statements. Different from them, the proposed method based on vision-language and visual contrastive learning can achieve superior results by using a more comprehensive text description. Table III and Table IV list the quantitative results on underwater and enhanced images datasets, respectively. The best and second-best results are highlighted in bold. HyperIQA, MUSIQ, CLIP-IQA exhibit significant performance degradation on the underwater dataset. This degradation may be due to the unique distortions in underwater images, which differ significantly from those in standard image quality assessments. The difference in performance is less in QADS, while the discrepancy in performance is least in GFIQA, which can be attributed to its sole composition of human facial images, thus presenting a limit range of content variability. This consistent outperformance on different types of IQA datasets underscores the robustness of SLIQUE and highlights the generalization of the representations learned by the model.
@@
(c) CONTRIQUE
(d) SLIQUE
Figure 4: 2D t-SNE visualizations of learned representations. For both CONTRIQUE and SLIQUE, we conducted 2D t-SNE visualization experiments using 7 different types of distorted images from 3 image databases and 3 different types of images from KADIS database.
Figure 5: 2D t-SNE visualizations of learned representations for 1,000 images in the Wild from KonIQ.
IV-C T-SNE Visualization of Quality Relevant Features
The success of SLIQUE relies on its image encoder’s ability to extract discriminative image features that can capture information related to factors affecting image quality such as semantic content, distortion characteristics and appearance properties. Fig. 4 shows examples of t-SNE visualization of the learned representations of SLIQUE and that of a recent method CONTRIQUE [31]. We used 750 images sampled from CSIQ (Synthetic distortions - #600) and User-Generated Content (UGC) images from KonIQ (UGC -#150) and pristine images from KADIS (Natural images - #150). It is clearly seen that, for the CONTRIQUE model, some features from differently distorted images are not well distinguished from each other, e.g., the JPEG and the Pink noise. In contrast, the representations learned by our new SLIQUE can better separate different distortion types which may explain why it has superior performances. Fig. 4 illustrates additional t-SNE visualizations of 300 pristine images from KADIS (100 each of buildings, cameras, and landscapes), distorted by Gaussian blur. The learned representations of CONTRIQUE model exhibit limited discriminative capacity across images with identical distortion types but varying content. In contrast, the learned representations of SLIQUE demonstrate more pronounced inter-class separation. Fig. 5 shows t-SNE visualizations of 1,000 KonIQ images, split evenly between low (<33) and high (>67) quality scores. The plot exhibits distinct clustering, with high-quality images (blue) and low-quality images (orange) forming separate groups. This pattern demonstrates the ability of SLIQUE to learn representations that effectively differentiate image quality. The overlapping region of the clusters represents instances where the predictions may diverge from ground truth. These results highlight the strong discriminative power of in distinguishing both content and quality variations, suggesting its effectiveness for image quality assessment tasks.
IV-D Cross Dataset Evaluations
To show the generalization ability of SLIQUE, we conduct cross dataset evaluations by training and testing on different datasets. Specifically, the ridge regression of SLIQUE is learned by all the data from a dataset while the parameters of image encoder $F$ are still frozen. Then, the model is tested on all the data from another dataset. Five datasets, including three synthetic and two authentic distortion datasets, are involved in the cross dataset evaluation. The results based on SROCC are presented in Table V, where the performances of two recently-developed models are also listed. As shown in Table V. The best and second-best performances are highlighted in bold. Comparing the results in Table II and Table V, we can see that cross-dataset evaluations are more challenging. It is mainly attributed to the difference of the data distribution between two IQA datasets, especially the difference between synthetic and authentic distortion datasets. Nevertheless, our model outperforms the Re-IQA and the SAWAR, which also take advantage of unsupervised learning to obtain quality-relevant features. It demonstrates that our joint contrastive learning is highly effective in representing the image quality.
TABLE V: Cross database evaluations of IQA models.
@@ IV-D Cross Dataset Evaluations
CLIVE CSIQ 0.680 0.656 0.721
CSIQ CLIVE 0.446 0.452 0.527
TID2013 CLIVE 0.423 0.411 0.562
CLIVE TID2013 0.436 0.452 0.531
@@ IV-E Ablation Studies
Individual contrastive learning. We first conducted an ablation study to validate the importance of joint contrastive learning by disabling one of the branches. When disabling the image-image branch, only image-language pairs are used to train the encoders $G$ and $F$ from scratch. Similarly, if the image-language branch is disabled, only image-image pairs are employed to train the image encoder $F$. The results based on SROCC are given in Table VI, which shows a performance reduction if we disable either branch. Moreover, we can see from the results that the visual contrastive learning is greatly important, and the vision-language contrastive learning also provides essential benefits.
TABLE VI: Performance of individual contrastive learning.
@@
Starting point of optimization. In the joint training of image-language branch and image-image branch, we optimize the language encoder $G$ and the image encoder $F$ from scratch. An alternative strategy is to optimize them (or one of them) from pre-trained models. Here we employed two well-known pre-trained models, i.e., CLIP and SIMCLR, as the starting point of optimization. The results on synthetic distortion datasets and authentic distortion datasets are recorded in Table VII and Table VIII, respectively. In the first row of each table, the language encoder $G$ is initialized by the pretrained CLIP weights, and the image encoder $F$ is initialized by the pretrained SIMCLR weights. Other cases are included in the following rows. As the SIMCLR only contains image encoder, it can not used to initialize $G$. From the results in Table VII and Table VIII, we find that the optimizations from pretrained models do not result in the best performances, although some performances from them are comparable with state of the art. It may be because that they are originally trained in object recognition or image classification tasks. Thus, using pretrained models is not optimal for the image quality assessment task.
TABLE VII: Optimization from pretrained models versus from scratch (synthetic distortion datasets).
@@
TABLE VIII: Optimization from pretrained models versus from scratch (authentic distortion datasets).
@@
Components in annotation texts. To investigate the impact of different components in the annotation texts on quality prediction, we conducted ablation studies by excluding some components in the textual descriptions, resulting in 4 cases. The results based on SROCC are shown in Table IX. In the first row of Table IX, we only keep the content annotation texts, and exclude both the distortion texts for synthetic distortion images and the appearance texts for authentic distortion images. The results in the first, second, and third rows respectively indicate that using only content text, distortion text, or appearance attribute text alone is insufficient. In the fourth row, we add the distortion information of synthetic distortion images to the textual descriptions. The results show that the performance on authentic distortion datasets KonIQ and CLIVE even decreases, although the performance on synthetic distortion datasets is improved. It may be attributed to the large differences between the synthetic distortion and the authentic distortion. Similarly, the results of adding the appearance texts are provided in the fifth row which is worse than the full model. In the sixth row, we report the results of excluding content texts, demonstrating that the content texts are also beneficial to the IQA task. The comparative analysis conducted above demonstrated the complementary relationship among all three annotation components, thus confirming their effectiveness and essential role in the overall process.
TABLE IX: Ablation studies on components in annotation texts. In each column, the top performance is boldfaced.
@@
TABLE X: SROCC performances of our model for different amount of training data.
@@
Usage of language feature. In our SLIQUE, only the image feature is employed in the prediction of visual quality of images. An alternative strategy is to use an image captioning method to produce a textual description for the image to be accessed. With such a strategy, the feature from language encoder can also be fed to the ridge regression. Here we employed the captioning method in [45] to produce annotation texts. The results on synthetic distortion datasets and authentic distortion datasets are recorded in Table XIII and Table XIV, respectively. From the results, we can see that using the language feature is not helpful. It is mainly because the texts produced from the captioning method contain little distortion and appearance information. Furthermore, even if the captioning method can produce perfect texts for the task, the language feature would be redundant due to its similarity to the corresponding image feature.
TABLE XI: Performance comparisons of SLIQUE and TCL retrained on Q-Pathway and TADAC on IQA databases with synthetic distortions.
Model LIVE CSIQ TID2013 KADID Weighted Average
[…]
TCL(Q-Pathway) 0.953 0.942 0.946 0.924 0.784 0.778 0.869 0.839 0.862 0.838
TCL(TADAC) 0.964 0.959 0.958 0.951 0.863 0.859 0.921 0.911 0.914 0.905
SLIQUE(Q-Pathway) 0.971 0.972 0.948 0.957 0.839 0.886 0.892 0.885 0.889 0.895
SLIQUE(TADAC) 0.982 0.982 0.966 0.973 0.884 0.899 0.957 0.959 0.943 0.948
TABLE XII: Performance comparisons of SLIQUE and TCL retrained on Q-Pathway and TADAC on IQA databases with authentic distortions.
Model KonIQ CLIVE FLIVE SPAQ Weighted Average
[…]
TCL(Q-Pathway) 0.672 0.707 0.674 0.706 0.503 0.533 0.834 0.829 0.526 0.556
TCL(TADAC) 0.811 0.831 0.708 0.749 0.563 0.589 0.858 0.852 0.593 0.618
SLIQUE(Q-Pathway) 0.891 0.906 0.795 0.819 0.674 0.685 0.916 0.920 0.723 0.739
SLIQUE(TADAC) 0.916 0.929 0.897 0.901 0.720 0.754 0.921 0.925 0.744 0.774
Benefits of TADAC dataset. To show the benefits of our TADAC dataset, we would like to re-train the text encoder $G$ and the image encoder $F$ by using another image-text dataset, known as Q-pathway [65]. Then, the regression is learned on a traditional IQA dataset. Finally, we record the model performances on the traditional dataset. Here, the 8 traditional datasets involved in Section IV.B are tested. The results based on synthetically distorted images are reported in the third row of Table VI, while the results on authentic distortion datasets are given in the third row of Table VII. By comparing the results in the last two rows of both tables, we can see that the performances based on the TADAC are better that those based on the Q-pathway. A probable reason is that the human feedbacks in the Q-pathway might be noisy due to different subjective scales of individuals with various cultural and educational backgrounds. In both tables, we further include the results from a competitive IQA model, known as TCL [25], which also uses image-text pairs to train an image encoder. We train the TCL encoder by using the TADAC and the Q-pathway, respectively. The performances of TCL are provided in the first two rows of Tables VI and VII, which also demonstrate the benefits of our TADAC dataset. Besides, as both results in the first and the third rows are based on the Q-pathway, we can conclude that the SLIQUE outperforms the TCL. The same conclusion can be drawn by comparing the results in the second and the last rows.
TABLE XIII: Exploring more features to use in evaluation of synthetic distortion datasets.
@@
TABLE XIV: Exploring more features used in the evaluation of authentic distortion datasets.